JUNIOR4256:

Brittany is making hair bows to sell at a craft show. Each hair bow requires34 yard of ribbon. Brittany plans to make 50 bows. How many yards of ribbon does Brittany need to purchase?

Convert:

34*50=1700 yards

jhonyy9:

so just use the rule of 3 simple and write 1 bow requires 34 yard 50 bows requires x and now cross multiplie to calcule the x value x = 50*34/1 = 50*34 = ?